OpenStudy (anonymous):

Find the derivative dy/dx of: y=(2x^3-4x+1)^-6 The answer is: -12(3x^2-2)(2x^3-4x+1)^-7 I'm able to go as far as: y=(2x^3-4x+1)^-6 u=(2x^3-4x+1) y=u^-6 dy/du= -6u^-7 du/dx= (6x^2-4) dy/dx=(dy/du)(du/dx) dy/dx (y)= -6(2x^3-4x+1)^-7(6x^2-4) I can't get beyond that, however. If you can enlighten me as to what I'm doing wrong, I'll let you be co-ruler when I take over the world.

OpenStudy (anonymous):

what makes you think you are doing something wrong? Looks fine.

OpenStudy (anonymous):

\[\frac{dy}{dx} = -6\frac{(6x^2-4)}{(2x^3-4x+1)^7} = -6\times2 \frac{(3x^2-2)}{(2x^3-4x+1)^7}=-12\frac{(3x^2-2)}{(2x^3-4x+1)^7}\]